Calculate the total pressure,in atm, in a tank of gas if it is composed of 205,333Pa N2, 950. mmHg O2 and 1.285 barN2O
1
Expert's answer
2020-12-16T07:47:37-0500

1Pa= 9.86923-6 atm

1bar=0.986923 atm

1mmHg=0.00131579atm

N2=205333*9.86923-6=2.0265atm

O2=950*0.00131579=1.2500atm

N2O=1.285*0.986923=1.2682atm

Total pressure in atm= (2.0265+1.2500+1.2682) atm

=4.5447atm